OpenStudy (shadyykatie):

if angles 1 and 2 are a linear pair, and m<1 = 3x + 12 and m<2 = x + 44, find the measures of these two angles.

OpenStudy (anonymous):

if two st lines intersect each other then he sum of the adjacent angles sum up to 180 and the vertically opposite angle so formed are equal

OpenStudy (shadyykatie):

linear pair means the sum of the angles is 180. (3x + 12) + (x + 44) = 180 solve x. I have trouble finishing it there!

OpenStudy (anonymous):

(3x + 12) + (x + 44) = 180 so 3x+x =180 -12 -44 4x=124 x=124/4 x=31

OpenStudy (hesan):